OpenStudy (anonymous):

if f(x)= log {(1+x)/(1-x)} then f (x) is a) an even function b) an odd function c) neither even nor odd d) a constant term

OpenStudy (jack12):

no im sorry

OpenStudy (jack12):

the answer is C

OpenStudy (anonymous):

i don't think so !

terenzreignz (terenzreignz):

LOL

ganeshie8 (ganeshie8):

evaluate f(-x)

terenzreignz (terenzreignz):

Someone has pretty powerful instincts :D

OpenStudy (jack12):

what do you think it is a constant term?

OpenStudy (jack12):

E its a cheese cake

terenzreignz (terenzreignz):

\[\Large f(x) = \log\left(\frac{1+x}{1-x}\right)\] Let's evaluate f(-x) \[\Large f(\color{red}{-x}) = \log\left(\frac{1-x}{1+x}\right)\]

OpenStudy (anonymous):

it is an odd function !

terenzreignz (terenzreignz):

Which, incidentally... is... \[\Large f(\color{red}{-x}) = \log\left(\frac{1+x}{1-x}\right)^{-1}\]

OpenStudy (jack12):

i didn't realize the denominator was inside the logarithm nvm

OpenStudy (jack12):

your tex wasn't formatted properly

ganeshie8 (ganeshie8):

you have powerful insticnts and you're superfast also krit ;p

OpenStudy (binary3i):

so @terenzreignz says f(-x)=-f(x)

OpenStudy (anonymous):

i have got another one !

ganeshie8 (ganeshie8):

okiee